OpenStudy (bruno102):

I don't understand this: An object is launched upward at 64 ft/sec from a platform that is 80 feet high. What is the objects maximum height if the equation of height (h) in terms of time (t) of the object is given by h(t) = -16t2 + 64t + 80?

OpenStudy (abb0t):

Find the derivative and find the roots of the function.

OpenStudy (bruno102):

I don't know how to do that.

OpenStudy (bruno102):

Actually, I believe I am understanding it.

OpenStudy (bruno102):

would it be 2, 1232? Or am I totally wrong

OpenStudy (phi):

h(t) = -16t2 + 64t + 80 is a parabola can you find its vertex? that will be when it is highest

zepdrix (zepdrix):

Hmm I'm not sure how to solve optimization problems like this without taking a derivative D: Hopefully phi can help.

zepdrix (zepdrix):

oh parabola :O hehe

OpenStudy (bruno102):

would it be 1232 feet?

zepdrix (zepdrix):

Factor -16 out of each of the first two terms, \[\large h(t)=-16(t^2-4t)+80\] Complete the square on the t's,\[\large h(t)=-16(t^2-4t+4)+64+80\]Which simplifies to,\[\Large h(t)=-16(t-2)^2+144\] Can you determine the vertex of the parabola now that it's in standard form?

OpenStudy (bruno102):

But what is t?

OpenStudy (anonymous):

the first coordianate of the vertex is always \(-\frac{b}{2a}\) which as you said, is 2 the second coordinate is \(h(2)\)

OpenStudy (anonymous):

i.e. plug in 2 where you see \(t\) to get the maximum height

OpenStudy (anonymous):

\[ h(2) = -16t^2 + 64t + 80\]\[ h(2) = -16\times (2)^2 + 64\times 2 + 80\]

OpenStudy (anonymous):

you do not need the derivative, and you do not need the roots you only need the magic formula for finding the first coordinate of the vertex of a quadratic \[y=ax^2+bx+c\] which is \(-\frac{b}{2a}\)

OpenStudy (anonymous):

you get \[h(2)=-16\times 4+128+80=64+80=144\]

OpenStudy (bruno102):

What are the numbers after the equal sign (64+80)?

OpenStudy (bruno102):

what happens to 128 in the equation if the answer is 144?

OpenStudy (anonymous):

are you asking how to compute \(h(2)\) ?

OpenStudy (bruno102):

I think so.

OpenStudy (anonymous):

replace every \(t\) in \(-16t^2+64t+80\) by a \(2\)

OpenStudy (bruno102):

Ok I did that and I get \[-16 \times 2^{2} +64 \times 2 + 80\] right?

OpenStudy (anonymous):

yes

OpenStudy (bruno102):

Ohhhh I get it now I was forgetting the negative sign of -64 when I added it to 128 and then 80. So the answer of the question is just 144.

OpenStudy (anonymous):

yes

OpenStudy (bruno102):

Ok thank you so much!

OpenStudy (anonymous):

yw